Grade 9Discuss with Askiitians Tutors

A body is moving in a circular path with acceleration a.if it`s speed gets doubled find the ratio of acceleration after and before the speed is changed

Karthik
8 Years agoGrade 9

1 Answer

Arun
8 Years ago

Centripetal acceleration=v^2/r=A

If velocity is doubled acceleration becomes 4v^2/r=4A

Ratio of initial and final acceleration=A/4A=1/4

 

hence ratio (after : initial) = 4:1

 

Regards

Arun (askIITians forum expert)